UT 20130123 Technical Observing

Observer: DMiller (LBTO) JHill (Tucson)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: DMiller (LBTO)
Telescope Support: KNewton (LBTO)
Instrument Support: TShih (LBTO)

Summary

First half of the night was TCS software testing for Build 2013A-test, DIMM and MCSPU. Since AGW2 "W" failed, we've kept the test TCS build for the second half of the night. (It will revert to 2012H in the morning.) Most of the second half was spent collecting field aberration data at LDG with the new collimation model.

Details

Testing the smarter acquisition algorithm

We spent an hour or so testing the new improved acquisition algorithm that now knows to choose the appropriate guide star even with a bright one off in the corner. (Issue 4379)

Corrected RFBG Pointing Model

04:xx Updated the RFBG pointing model to account for the DX M3 tip-tilt we did last night. RFBG_AGw_20111208.ptmod becomes RFBG_AGw_20130123.ptmod with IE=-43 CA=-124

Binocular Acquisitions

We tested the binocular acquistion algorithm (Issue 4328). We (MDD) also patched GCS to avoid the binocular timing delay when restoring guiding (Issue 4434).

DX M2 astigmatism

04:45 EL=65 DX Z6=-2534

04:50 JCG turns off the DX M2 astigmatism LUT.

04:57 EL=68 DX Z6=-1314 (present DX LUT apparently has the wrong sign!)

05:11 EL=33.5 DX Z6=-11256

05:28 EL=83.3 DX Z6=+1007

05:35 EL=50:17 DX Z6=-7038

To zenith to switch from MODS to LBC (after finding LFBG rotator locked out).

06:30 EL=78:42 DX Z6=-3571

LBC-Blue mixed with RFBG

SYNC presets from IRAF (RFBG) and LBC-Blue

LBC Guiding Glitch

06:45 near the time of a RFBG offset, there was a 1/3 arcsec guiding jump on LBC-Blue. We need to investigate this. Repeating the offset ~5 min later did not produce a glitch. Later investigation by MDD showed that the glitch was not correlated to the time of the offset.

MODS Image Quality

07:30 Reconfigure to MODS complete. Steve send preset and collimate. Seeing 1.0-1.2. No tails on guide image

07:45 Preset to several off-axis guide stars to see if vignetting of wfsc image cause tails. Guide probe fails to move to requested position several times the I would think would be reachable. John suggest we re-home the MODS guider stages. Apparently that worked.

08:05 John installs Dave A's new pointing and collimation models

We discovered that PSF has a line length limit. You can't use 10 digits for the Temp1 Temp2 line.

The collimation model "Ashby2" aka "Updated_Collimation_and_Pointing_Sep2012_plus_LT" gave poor pointing results (arcminute errors).

08:39 We switched to "Ashby3" collimation model, and the previous MODS pointing model.

MODS pupil shift parameters

Best fit to 5 nights of on-sky data using new IDL routine to fit outer edge of pupil

X Pupil Shift y = 379.23 + 0.14286 * x

Y Pupil Shift
y = 247.8 - 0.14756 * x

Thus, WFS_hotspot_x, y = 379.23, 247.8

WFS_r_to_pupil_shift = 0.145

Installed in MODS_L.cfg

MODS Field Data at 60 Elevation

This is using the new collimation model "Ashby3" aka "Update_Collimation_NO_Pointing_Sept2012_plus_LT"

08:55 field_collect, BS9126_field.list
  • FWHM = 1.1
  • directory /home/aoeng/data/20130123/el_60
 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -22.8", Delta IE =  22.8"
         Tilt = -18.9", Delta CA =  18.9"

MODS Field Data at 66 Elevation

10:33 Apply -40 arcsec TIP pointing correction to SX M2. dIE=+40
  • directory /home/aoeng/data/20130123/tip-40

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -45.1", Delta IE =  45.1"
         Tilt =   5.1", Delta CA =  -5.1"

MODS Field Data at 71 Elevation

11:00 Apply +40 arcsec TIP pointing correction to SX M2. dIE=-40
  • directory /home/aoeng/data/20130123/tip+40

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  45.5", Delta IE = -45.5"
         Tilt =   1.9", Delta CA =  -1.9"

MODS Field Data at 75 Elevation

11:35 Proper high elevation field data set
  • directory /home/aoeng/data/20130123/el_75

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -11.9", Delta IE =  11.9"
         Tilt =  20.9", Delta CA = -20.9"

MODS Field Data at 77 Elevation

12:00 Apply -40 arcsec TILT pointing correction to SX M2. dCA=+40
  • directory /home/aoeng/data/20130123/tilt-40

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  37.7", Delta IE = -37.7"
         Tilt =  83.3", Delta CA = -83.3"

MODS Field Data at 78 Elevation

12:30 Apply +40 arcsec TILT pointing correction to SX M2. dCA=-40
  • directory /home/aoeng/data/20130123/tilt+40
 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   6.7", Delta IE =  -6.7"
         Tilt = -21.0", Delta CA =  21.0"

Testing StepFocus command

12:56 We sent several StepFocus commands to LDG to verify that the zero amplitude command updates the mirror positions w/o pausing guiding. This worked correctly as advertised in 2013A-test. We also sent non-zero amplitudes to verify that pausing still works in that instance.

High Order Zernikes Z12-Z22

13:00 Were tested here.

Open Loop Tracking

13:09:20 Stopping sending guide corrections to test open-loop tracking at LDG (no special pointing correction). AZ=323 EL=75 This is the old pointing model on the new collimation model, so not a very well fit model. Collimation is not being updated at all. Initial drift was 1 arcsec in 5 minutes.

This test ended at 13:25:06.

Close

13:26 Close for impending dawn.

-- DougMiller - 23 Jan 2013
Topic revision: r7 - 28 Jan 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback